OpenStudy (anonymous):

consider a hyperbola with vertices (5, 0) and (–5, 0) and asymptotes y = ± 8/5 x. Write an equation for the hyperbola.

OpenStudy (mertsj):

Whoops. Put a minus sign between the two terms. That is an ellipse.

OpenStudy (mertsj):

\[\frac{x^2}{25}-\frac{y^2}{64}=1\]